ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30B50 Fedynyak, Stepan, Lviv State University, Ukraine Derivative of a Dirichlet series For an analytic in {s : Res < A}, A ∈ (−∞, +∞], function F, represented by an absolutely convergent Dirichlet series with exponents growing to +∞ and satisfies the condition _{σ→A} \frac{ln M(σ,F)}{Φ(σ)} = 1, where M(σ, F) = sup\{|F (σ + it)| : t ∈ ℝ} and Φ be a convex on (−∞, A) function. The asymptotic behavior of the function S_{1}(σ, F) = \frac{M(σ,F′)}{M(σ,F)} is studied.